# Fairness-in-the-Workflow: How Machine Learning Practitioners at Big Tech Companies Approach Fairness in Recommender Systems

### v2

@@ 5.1. RS Practitioners Struggle to Define Fairness in the Context of Recommender Systems
Defining Fairness.
Practitioners attempted to assess and improve fairness within two general categories. The first category is fairness as mandated by existing rules or regulations. For example, an RS that discriminated against job candidates from particular races would be considered unfair under US anti-discrimination law. The second category, encompassing all contexts not covered by existing rules or regulations, is perceived as more difficult and nuanced. One common frustration was that practitioners only have a vague goal (for example, not exposing content creators to accounts that might harass them), but are not armed with specific tools in their organization to reach the goal—or even a standardized method with which to approach the problem.
Adapting Pre-Existing Fairness Metrics and Navigating Documentation Debt.
While approaches for assessing and improving fairness have been proposed by fair ML literature (Chen et al., 2023), practitioners found them hard to directly adapt and ill-suited for their existing workflows. For example, P9 reported, “Even though we have a list of categories [over which to assess fairness] as you mentioned in research papers, we need to come up with a clear formula and implement it ourselves. It is really domain-specific.” Practitioners also struggled to build and re-use institutional knowledge about fairness in RS. This was due to the fact that RS practitioners faced documentation debt (Bender et al., 2021): prior fairness-related work at their organizations was not typically well-documented for future use. Half of practitioners interviewed therefore faced difficulties in re-using internal metrics that had been previously developed for fairness purposes. P7 shared, “For sure we can find their code, and comments, but without some context archived, for example, some documents, we are not sure whether we can use [them] directly.” Practitioners also felt ill-equipped to document their own fairness processes. As P3 stated, “Documentation is a good idea…however, I am not sure how specific of my debiasing should be included in the documentation.”
@@ 5.2. RS Practitioners Must Account for Multiple Types of Fairness Across Many Groups
Prior research emphasizes the dual role of recommender systems in simultaneously connecting and serving multiple stakeholders, including providers and end-users (Burke, 2017; Li et al., 2021b). This multifaceted interaction introduces complex challenges, especially in the context of fairness. The practitioners that we interviewed commonly made tradeoffs, often prioritizing one user group over others. However, deciding how to make these tradeoffs was a challenge, as highlighted by P9: “Tradeoffs of a different group of users, or accuracy of the recommendations are hard to be decided,” and P1, who pointed out a lack of a “compelling [Standard Operating Procedure] for those decisions.” Making tradeoffs, while common, may not always be necessary. Recent academic findings suggest that multi-objective optimization might not always adversely affect particular user groups (Ekstrand et al., 2022), indicating a disconnect between industry practices and academic research.
Prioritizing Between Different (Non-Conflicting) Types of Fairness.
@@
Practitioners felt that there was more of a “what to prioritize” question in RS compared with traditional predictive models. In a traditional ML pipeline, fairness assessment focuses on one particular target (e.g., demographic parity). However, the practitioners that we interviewed expressed that they have to balance multiple fairness-related considerations. For example, one participant needed to consider demographic parity for job seekers in one region—and simultaneously, ensure that advertisers across multiple regions were adhering to different sets of fairness-related regulatory requirements. Importantly, these fairness considerations were not necessarily in conflict with one another. However, because practitioners had limited time and resources available, they nevertheless needed to choose certain fairness ideals to prioritize before turning to others. P3 pointed out: “Debiasing for multiple purposes is crucial in ensuring a fair and unbiased platform for all stakeholders. It’s not just about the debiasing of content creators based on region, gender, and religion…but also about addressing other factors that can impact the platform…For instance, we need to make sure that products…are supported fairly [e.g., that products that are artificially cheap because of government subsidies designed to gain market dominance are downranked on our RS platform].”
@@
A significant challenge identified by practitioners involves ensuring fairness across a wide range of groups within RS platforms. Current metrics often fail to address the complexities arising from the large number of groups involved (and the ensuing intersectional groups). P2 described this challenge: “It’s hard to manage disparities between the numerous groups. With thousands of groups to consider, it becomes nearly impossible to ensure uniformity across all.” This sentiment was echoed by P9, who noted, “Most metrics are designed with the assumption of a few groups…They don’t cater to intra-group-level debiasing, making our adaptation efforts cumbersome.” Additionally, P7 highlighted the computational difficulties in handling sparse datasets resulting from subgroup division: “Most of the training data is too sparse, and it is hard for us to enforce the disparities between groups. I cannot look through thousands of groups to make sure that holds.”
@@ 5.3. Fairness Considerations are Complicated by Dynamic RS Environments
Developing Ad Hoc Metrics to Detect Unexpected Instances of Unfairness.
Practitioners engaged in building RS often create bespoke metrics in collaboration with their teams. P2 and P3 highlighted the personalized nature of these metrics, with P2 noting, “We proposed our own definitions of bias,” and P3 adding, “Most of the metrics we developed are intuition-based.” Such approaches underscore the reliance on practitioner expertise and intuition in assessing and improving fairness, as echoed by P6: “In general, I would like to write some hand-crafted codes to validate my hypothesis. Experience is a key to debiasing sometimes,” despite the availability of internal fairness tools. Nevertheless, practitioners acknowledged that their initial metrics often fell short in detecting unfairness, and they therefore engaged in iterative refinement of those metrics based on user feedback:
[P1]: It is not an easy job…so we need to start from somewhere. After doing some debiasing based on your experience, we look forward to getting feedback from real-world users to further debiasing.
[P7]: It sounds a bit contradictory. We have a model first and keep debiasing the model based on our user’s new data. It although risks bringing in some confirmatory bias, it works for now.
@@
Practitioners also shed light on the challenges of assessing and improving fairness in online (deployed), as compared to offline, RS. This challenge is particularly acute in RS given the volume of data, the subtleties of human interactions, and the potential for user feedback loops to amplift latent sources of unfairness. P1 and P2 emphasized the difficulty in uncovering unexpected fairness issues in deployed systems, with P1 stating, “As the bias definitions are so different, and there are so many different user groups, those biases are hiding well behind the recommender systems that we built,” and P2 reflecting, “Most of the time, we have to look carefully to find those hidden biases. It is not fun.” Despite finding academic metrics less directly applicable (see §5.1), practitioners acknowledged the value of guidelines from research in providing a foundation and confidence in their decision-making processes. For example, P7 stated, “Though sometimes [checklists] are not directly related, I can still check those and they help me think through my reasoning. It made me somehow more confident in my decision.”
@@ 6. Organizational Challenges Arising Across Teams When Incorporating Fairness
In this section, we describe the key organizational challenges faced by RS practitioners who seek to incorporate fairness considerations into their workflows. As with the technical challenges outlined in §5, these findings emerge out of the challenge-centric thematic analysis described in §3. Table 2 details the organizational challenges most commonly reported by practitioners.
@@ 6.1. RS Practitioners Lack Time to Engage in Collaborative Fairness Work
Practitioners universally described their workflows as highly iterative and time-sensitive. As a result, most felt they had a very limited amount of time to focus on fairness, either internally or in collaborations with the fairness team. Nine reported that they spent roughly 5% percent of their work time on fairness-related tasks, and two said they spent up to 10% of their time. As P7 put it, “The negative effect spurred from the platform I worked for before is not so ‘bad.’ We don’t have the incentive to invest half of our time…to comprehensively debias.” When prompted to reflect on this allocation, most practitioners indicated that, while allocating more time for fairness-related tasks could be helpful, they were neutral about doing so. P1 reported, “Even though I wish I could have more time on this, I understand why this proportional[ly] makes sense. From the perspective of our org, prioritizing efforts on improving ‘important’ metrics brings more impact.” In other words, fairness issues did not rise to the level of a disruptive emergency, and thus were not prioritzed by management in evaluating project outcomes.
@@ 6.2. RS Practitioners Lack a ‘Fairness Lingua Franca’ to Facilitate Cross-Team Communication
Engaging with external fairness teams revealed a nuanced challenge: attaining mutual understanding and converging perspectives. Practitioners valued the fairness teams’ feedback but sometimes struggled with its practical integration due to a few key complexities. Firstly, the inherent structure of RS as complex deep learning models sometimes made feedback from fairness teams difficult to implement. P9 observed, “sometimes…[a] challenge is fundamentally unsolvable, rooted in the deep learning models.” P1 shared, “Their feedback could be something I agree with, but I cannot solve.” Secondly, the multidisciplinary nature of fairness teams led them to use terminology that was not readily understood by technical teams, affecting inter-team communication. P9 commented on the influence of varied backgrounds and perspectives on understanding: “sometimes I don’t understand some phrases they used in their feedback, for example, ‘inductive bias from recommender systems,’ and need to ask for more explanations.” P6 added, “Some of my colleagues are from the field of psychology and communication, and some of those terms I didn’t know before.” Thirdly, practitioners desired clearer guidance from fairness teams, particularly regarding model development. P7 said, “Once they mentioned a certain aspect of my model might introduce bias, I wished to understand the reasoning.” Finally, fairness teams were often involved only after models were deployed for online testing. P9 suggested, “Early involvement of the fairness team would have facilitated a closer working relationship.” P2 noted, “Sometimes there’s a lag in syncing up; feedback might refer to previous versions of our recommender system designs.”
@@
These challenges starkly contrast with how RS practitioners interact with legal teams. Due to the sensitive nature of end-user data, practitioners regularly interact with their legal team to confirm the scope of data that can be gathered. With increasing surveillance from authorities and global concerns over privacy on RS, one might expect the relationship between practitioners and internal policymakers to be fraught. However, practitioners reported that communication with legal teams was clear and direct. Per P2, “If they said ‘no,’ then it is a clear ‘no,’ and we know the reasons and risks behind their decision.” Similarly, P3 stated, “Their guidelines are very clear, particularly when compared with fairness guidelines, which we don’t really have.” Clear and open guidelines from legal teams made it easier for practitioners to work with them, and most practitioners acknowledged the role of the legal team in avoiding risks. For instance, P2 mentioned that the “legal team serves as the gatekeeper to keep us away from legal risks, and help us comply well with our company’s vision and values.” P1 summarized the issue succinctly: “Our collaborations [with fairness teams] could be streamlined if there were guidelines as clear as those from legal teams.”

### v3

@@ 5.1. RS Practitioners Struggle to Define Fairness in the Context of Recommender Systems
Defining Fairness
RS practitioners attempt to assess and improve fairness within two general categories. The first category is fairness as mandated by existing rules or regulations. For example, an RS that discriminates against job candidates of particular races would be considered unfair under US anti-discrimination law. Even when participants had a legally mandated fairness goal in mind, they reported struggling with how to achieve it. Participants felt they were not armed with specific tools or standardized methods to reach even well-specified fairness goals. This echoes Smith et al. (2023a), who found that as RS practitioners tried to select appropriate fairness metrics, they struggled with not only normative questions like which stakeholders should be prioritized, but also practical ones like whether to measure fairness across an entire distribution of recommendations or only across a top-k set of rankings.
Participants perceived the second category, encompassing all contexts not covered by existing rules or regulations, as more difficult and nuanced. This category includes cases where attributes that are legally protected in some contexts are not legally protected in others. For example, although racial discrimination is illegal in employment RS, it is not illegal in dating RS (Hutson et al., 2018). Thus, it is not clear how consistently to apply the same fairness metrics and techniques across contexts. This category also includes fairness across attributes that are not legally protected in any context, such as the popularity of a provider’s content. These attributes frequently conflict and can compound to the point of being overwhelming. For example, P7 stated that, “Each module of my recommender system requires different debiasing,” with goals ranging from avoiding a “rich gets richer” feedback loop to providing recommendations of equal quality to different groups of end users. This question of what to do in the absence of legal requirements is one that industry AI teams consistently grapple with (Quiñonero Candela et al., 2023).
Adapting Pre-Existing Fairness Metrics and Navigating Documentation Debt
While approaches for assessing and improving fairness in RS have been proposed by researchers (see §2.1), participants found them hard to directly adapt and ill-suited for their existing workflows. In some cases, participants simply did not have the time to find existing metrics:
P4: I didn’t find very useful literature but honestly, I didn’t have much time to do a serious literature review and read papers either.
Participants also struggled to build and re-use institutional knowledge about fairness in RS. This was due to the fact that they faced documentation debt (Bender et al., 2021): prior fairness-related work at their organizations was not typically well-documented for future use. Half of practitioners interviewed therefore faced difficulties in re-using internal metrics that had been previously developed for fairness purposes:
P7: For sure we can find [old] code and comments, but without some context archived, some documents, we are not sure whether we can use [them] directly.
Participants also felt ill-equipped to document their own fairness processes. As P3 stated, “Documentation is a good idea…however, I am not sure how specific [I should be].” These challenges strongly echo those surfaced by prior work on NLP, ML, and AI practitioners (Harvey et al., 2025; Holstein et al., 2019; Ojewale et al., 2025; Lee and Singh, 2021).
@@ 5.2. RS Practitioners Must Account for Multiple Types of Fairness Across Many Groups
RS simultaneously connect and serve multiple stakeholders, including providers and end-users (Burke, 2017; Li et al., 2021b; Smith et al., 2023a). This multifaceted interaction introduces complex challenges, especially in the context of fairness. Participants commonly made tradeoffs, often prioritizing one user group over others. However, deciding how to make these tradeoffs was a challenge, as highlighted by P9, who felt that, “Tradeoffs [between] groups of users…are hard to [decide],” and P1, who pointed out a lack of a “compelling [standard operating procedure] for those decisions.” Deciding which stakeholders to prioritize is a known challenge across fair ML (Madaio et al., 2022). However, in the context of RS, it may not always be necessary. Recent academic findings suggest that multi-objective optimization might not always adversely affect particular user groups, indicating a disconnect between industry practices and academic research in the RS domain (Ekstrand et al., 2022).
Prioritizing Between Different (Non-Conflicting) Types of Fairness
@@
Participants felt that there was more of a question of what to prioritize in the context of RS as compared to predictive models. In a predictive ML pipeline, fairness assessment typically focuses on one particular target (e.g., demographic parity). However, the practitioners that we interviewed expressed that they had to balance multiple fairness-related considerations. For example, one participant needed to consider demographic parity for job seekers in one region—and simultaneously ensure that advertisers across multiple regions were adhering to different sets of fairness-related regulatory requirements. Importantly, these fairness considerations were not necessarily in conflict with one another. However, because practitioners have limited time and resources available, they nevertheless need to choose certain fairness ideals to prioritize before turning to others:
P3: Debiasing for multiple purposes is crucial in ensuring a fair and unbiased platform for all stakeholders. It’s not just about the debiasing of content creators based on region, gender, and religion…but also about addressing other factors that can impact the platform…For instance, we need to make sure that products…are supported fairly [e.g., that products that are artificially cheap because of government subsidies designed to gain market dominance are downranked on our RS platform].
Because most work in fair ML focuses on prediction, in which different types of fairness are often incompatible with one another, prioritizing between different non-conflicting types of fairness represents a distinct challenge to doing fair ML in practice. While prior work has found that practitioners struggle to make value judgments about what to prioritize when faced with capacity constraints, that work tends to focus on the relationship between fairness and other, non-conflicting but not necessarily fairness-related, desiderata. For example, fairness efforts may be de-prioritized due to a focus on user privacy (Holstein et al., 2019) or an organizational emphasis on maximizing user engagement or product functionality (Rakova et al., 2021). Other work rightly notes that different fairness paradigms (e.g., procedural fairness and outcome-based fairness) do not necessarily conflict with one another and thus can be simultaneously satisfied by combining procedural fairness interventions (e.g., checklists) with technical ones intended to improve an outcome-based fairness metric (Madaio et al., 2020). In contrast, we focus on the idea that even within a single paradigm (e.g., outcome-based fairness), there are multiple non-conflicting fairness definitions that RS systems could simultaneously satisfy. Nevertheless, we find that even when a given practitioner is tasked with improving fairness (i.e., when fairness in general is not de-prioritized), that practitioner must essentially rank their planned fairness interventions (including both procedural and outcome-focused interventions) and prioritize accordingly.
@@
A significant challenge identified by participants involves ensuring fairness across a wide range of groups within RS platforms. Current metrics often fail to address the complexities arising from the large number of groups involved (and the ensuing intersectional groups). Multiple participants raised this challenge:
P2: It’s hard to manage disparities between numerous groups. With thousands of groups to consider, it becomes nearly impossible to ensure uniformity across all.
P7: Most of the training data is too sparse, and it is hard for us to enforce [fairness] between groups. I cannot look through thousands of groups to make sure [fairness] holds.
Participants also struggled to find metrics that went beyond group fairness. For example, P9 noted that “Most metrics are designed with the assumption of a few groups…They don’t cater to intra-group-level debiasing, making our adaptation efforts cumbersome.” This challenge is exacerbated not only by the large scale of RS platforms, but also by the previously mentioned challenges of determining how to define and prioritize fairness metrics that are not required by law. In predictive ML settings, fairness assessments often take cues from anti-discrimination law by assessing differences in outcomes across legally protected demographic groups. In contrast, the RS practitioners we interviewed appear to rely less on anti-discrimination law as a blueprint for fairness assessments.
@@ 5.3. Fairness Considerations are Complicated by Dynamic RS Environments
Developing Ad Hoc Metrics to Detect Unexpected Instances of Unfairness
Practitioners engaged in building RS often create bespoke metrics in collaboration with their teams. Such approaches underscore the reliance on practitioner expertise and intuition in assessing and improving fairness, as highlighted by P6: “In general, I would like to write hand-crafted code to validate my hypothesis—experience is key to debiasing.” Nevertheless, participants acknowledged that their initial metrics often fell short in detecting unfairness, and they therefore engaged in iterative refinement of those metrics based on user feedback:
P1: It is not an easy job…so we need to start from somewhere. After doing some debiasing based on [our] experience, we look forward to getting feedback from real-world users…We do everything that we can do. However, it is not rare we detect some other bias after the model is [deployed].
P7: It sounds a bit contradictory. We have a model first and keep debiasing the model based on our users’ new data. Although [it] risks bringing in some confirmatory bias, it works for now.
Again, this highlights a challenge specific to measuring and improving fairness in RS: offline testing is inherently limited because it requires detailed predictions of how users will interact with deployed RS. This means that RS practitioners must either develop models of predicted user behavior or must wait until systems are deployed for comprehensive testing. This stands in stark contrast to most predictive ML systems, which are often testable using historical data, and even to generative AI systems, which may be tested through red teaming.
@@
Practitioners also face the challenge of assessing and improving fairness in deployed (as opposed to pre-deployment) RS. This challenge is particularly acute in RS given the volume of data, the subtleties of human interactions, and the potential for user feedback loops to amplify latent sources of unfairness. P1 emphasized the difficulty in uncovering unexpected fairness issues in deployed systems, stating, “As the bias definitions are so different, and there are so many different user groups, those biases are hiding behind the systems that we built.” Despite finding academic metrics less directly applicable (see §5.1), participants acknowledged the value of research in providing a foundation for their decision-making processes. For example, P7 stated, “Though sometimes [checklists] are not directly related, I can still check [them] and they help me think through my reasoning. It made me more confident in my decision.” At the same time, however, participants did not feel comfortable using pre-defined unfairness mitigation techniques (i.e., from academic research) on deployed systems. As P2 put it, “There is unknown risk of using those approaches. If the model performance drops a lot, it is a disaster.” This challenge is exacerbated by the fact that there is a comparatively small amount of fair ML research focused on post-deployment fairness improvements (Black et al., 2023).
@@ 6. Organizational Challenges Arising Across Teams When Incorporating Fairness
In this section, we describe the key organizational challenges faced by RS practitioners who seek to incorporate fairness considerations into their workflows (summarized in Table 2). The challenges we identify—namely, a lack of time to engage in fairness work and poor cross-team communication—echo those identified by prior interview studies conducted on AI and ML practitioners (Harvey et al., 2025; Deng et al., 2023b; Rakova et al., 2021; Balayn et al., 2023). However, we also identify a notable exception to the challenge of cross-team communication: while the RS practitioners we interviewed struggled to communicate with fairness teams, they overwhelmingly described clear and productive communications with legal teams. Thus, in this section, we contrast between the two to offer new insight into why RS practitioners struggle to develop a “fairness lingua franca.”
@@ 6.1. RS Practitioners Lack Time to Engage in Collaborative Fairness Work
Participants universally described their workflows as highly iterative and time-sensitive. As a result, most felt they had a very limited amount of time to focus on fairness, either internally or in collaborations with the fairness team: all reported spending 10% or less of their time on fairness-related tasks. As P7 put it, “The negative effect spurred from the platform I worked for before is not so ‘bad.’ We don’t have the incentive to invest half of our time…to comprehensively debias.” When prompted to reflect on this allocation, most participants indicated that, while allocating more time for fairness-related tasks could be helpful, they were neutral about doing so:
P1: Even though I wish I could have more time on this, I understand why this proportionally makes sense. From the perspective of our [organization], prioritizing efforts on improving ‘important’ metrics brings more impact.
In other words, fairness issues did not rise to the level of a disruptive emergency, and thus were not prioritzed by management in evaluating project outcomes. This challenge is not unique to RS, and has been identified by researchers time and time again as a particular barrier to doing fairness work in practice (Veale et al., 2018; Holstein et al., 2019; Madaio et al., 2020; Rakova et al., 2021; Madaio et al., 2022; Deng et al., 2023b; Balayn et al., 2023; Harvey et al., 2025).
@@ 6.2. RS Practitioners Lack a ‘Fairness Lingua Franca’ to Facilitate Cross-Team Communication
Engaging with external fairness teams revealed a nuanced challenge: attaining mutual understanding and converging perspectives. Participants valued fairness teams’ feedback but sometimes struggled to integrate it due to a few key complexities. First, the complexity of RS sometimes made feedback from fairness teams difficult to implement, as highlighted by P1: “Their feedback could be something I agree with, but I cannot solve,” and by P9: “Sometimes…[a] challenge is fundamentally unsolvable, rooted in the deep learning models.” Second, the multidisciplinary nature of fairness teams led them to use terminology that was not readily understood by technical teams, affecting inter-team communication:
P6: Some of my colleagues are from the field of psychology and communication, and some of those terms I didn’t know before.
P9: Sometimes I don’t understand phrases they use in their feedback, for example, ‘inductive bias from recommender systems,’ and need to ask for more explanations.
Third, participants desired clearer guidance from fairness teams, particularly regarding model development. For example, P7 explained that, “Once [the fairness team] mentioned a certain aspect of my model might introduce bias, I wished to understand the reasoning.” Finally, fairness teams were often involved only after models were deployed for online testing. As P9 put it, “I could work with [fairness] closer if they could have joined [our project] earlier.”
We are not the first to identify communication challenges as a barrier to assessing and improving fairness in practice: notably, Deng et al. (2023b) interviewed AI practitioners to understand “cross-functional collaboration” in industry. They found that practitioners addressed communication challenges by taking on “bridging roles,” in which they championed fairness efforts by initiating cross-team meetings and developing educational resources for their colleagues. While beneficial to their teams, practitioners who took on bridging roles struggled with under-recognition and unreasonable expectations, as their existing full-time jobs did not typically include explicit fairness-related responsibilities (Deng et al., 2023b). Interestingly, even though the RS practitioners we interviewed all worked with dedicated fairness teams—theoretically reducing the need for informal bridging roles—they still struggled with cross-team communication.
@@
Importantly, however, cross-team communication challenges were not ubiquitous: participants overwhelmingly drew a sharp contrast between their interactions with fairness teams and their interactions with legal teams. Due to the sensitive nature of end-user data, practitioners regularly interact with their legal team to confirm the scope of data that can be gathered. With increasing surveillance from authorities and global concerns over privacy in RS, one might expect the relationship between practitioners and internal policymakers to be fraught. However, participants reported that communication with legal teams was clear and direct. Per P2, “If they said ‘no,’ then it is a clear ‘no,’ and we know the reasons and risks behind their decision.” Similarly, P3 stated, “Their guidelines are very clear, particularly when compared with fairness guidelines, which we don’t really have.” Clear and open guidelines from legal teams made it easier for practitioners to work with them, and most participants acknowledged the role of the legal team in avoiding risks. P1 summarized the issue succinctly: “Our collaborations [with fairness teams] could be streamlined if there were guidelines as clear as those from legal teams.” Another thing that the technical practitioners we interviewed appreciated about legal teams was their early engagement—legal teams become involved in fairness work during the offline prototyping phase, while fairness teams typically do not engage until the online collaborative phase. Friction in cross-team understanding was a hindrance:
P9: Sometimes, [the fairness team asks] for a lot of abbreviations…which can be found in our review documents. Personally, I could understand that joining in the middle of a project, it is hard for them to know a lot of discussions and terms that we used.
Similarly, P2 felt that fairness teams did not always provide actionable recommendations, saying “[Fairness is] sometimes not on the same page—they gave feedback based on outdated [RS] designs.” These observations suggest potential paths forward for improving communication between fairness and technical teams.